# Diagnosing and Mitigating Modality Interference in Multimodal Large Language Models

### v2

@@ 4 Methods
To mitigate modality interference and enhance cross-modality competency, we propose a unified perturbation-aware training framework that introduces interventions at both the input level (on $X_{I}$ and $X_{T}$) and the representation level (on $Z_{I}$ and $Z_{T}$) with consistency regularization. Overall pipeline is displayed in  Figure˜3.
@@ 4.1 Perturbation-based Data Augmentation
To increase causal robustness along the desired paths $X_{I}\rightarrow Z_{I}\rightarrow A$ (for image-heavy tasks) or $X_{T}\rightarrow Z_{T}\rightarrow A$ (for text-heavy tasks), we first adopt a diverse and causally grounded data augmentation method by augmenting each sample with both heuristic perturbations and training time adversarial perturbations.
@@
We sample $N_{\text{img}}$ and $N_{\text{text}}$ examples from $\mathcal{D}^{\text{img}}$ and $\mathcal{D}^{\text{text}}$ to construct $\mathcal{B}^{\text{img}}_{\text{orig}}$ and $\mathcal{B}^{\text{text}}_{\text{orig}}$ respectively, and the remaining $N_{\text{vqa}}$ examples are VQA samples from $\mathcal{D}^{\text{vqa}}$ to construct $\mathcal{B}^{\text{vqa}}$ . With dynamically generated perturbed variants for each image-heavy and text-heavy sample, the final training batch is:
@@
This objective enables full supervised fine-tuning across modality-specific samples with their heuristic perturbed variants and VQA samples, encouraging the model to jointly learn robust unimodal reasoning and generalizable multimodal alignment.
@@
While heuristic perturbations simulate realistic but limited modality noise at the input level, they may not fully capture the worst-case failure modes of MLLMs, especially under complex spurious alignments in the representation space. To overcome this limitation, we introduce a stronger and more generalizable intervention through adversarial training. These perturbations simulate worst-case alignment disruptions during training, serving as targeted interventions on latent nodes $(Z_{I},Z_{T})$ to reveal the Direct Causal Effect (DCE) of irrelevant modalities on $A$. By optimizing the model under such adversarial conditions, we reduce the model’s reliance on spurious cross-modal signals and reinforce task-relevant causal pathways. Specifically, we apply Projected Gradient Descent (PGD) DBLP:conf/iclr/MadryMSTV18 ; ptp at training time to perturb the multimodal token embeddings $(Z_{I},Z_{T})$ directly. It constructs perturbations $\delta=(\delta_{I},\delta_{T})$ in the latent space that maximize the model’s predictive loss:
@@
where $\epsilon$ bounds the perturbation strength and $f$ denotes the model’s prediction function. We implement this objective via $n$-step iterations, where at each step $t$, the perturbation is updated as:
$\delta^{(t+1)}=\Pi_{\|\delta\|_{\infty}\leq\epsilon}\left(\delta^{(t)}+\alpha\cdot\left(\nabla_{\delta}\mathcal{L_{\text{cls}}}(f(Z+\delta^{(t)}))\right)\right),$ (5)
where $\Pi$ is the projection operator onto the $\ell_{\infty}$-ball, $\alpha$ is the step size, and $Z=[Z_{I};Z_{T}]$ is the concatenated embedding. By injecting these adversarial perturbations into $(Z_{I},Z_{T})$ during training, we simulate worst-case interventions on the model’s internal decision pathways, exposing and correcting spurious dependencies between irrelevant modality features and prediction $A$. In practice, we restrict the perturbation to the task-irrelevant modality using a modality-specific binary mask $M\in\{0,1\}^{L\times d}$, where $L$ is the sequence length and $d$ is the dimension. Let $\mathbf{E}\in\mathbb{R}^{L\times d}$ denote the multimodal token embedding sequence, where projected visual features are already interleaved into the textual token sequence llava ; qwen25vl ; internvl3 . We construct adversarially perturbed embeddings $\tilde{\mathbf{E}}$ as:
@@
where $\delta\in\mathbb{R}^{L\times d}$ is the perturbation and $\odot$ denotes element-wise masking. We then optimize $\delta$ via $T$ steps of projected gradient ascent:
$\delta^{(t+1)}=\Pi_{\|\delta\|_{\infty}\leq\epsilon}\left(\delta^{(t)}+\alpha\cdot\left(\nabla_{\delta}\mathcal{L}(f(\mathbf{E}+\delta^{(t)}\odot M))\right)\right),$ (7)
where $\alpha$ is the step size and $\Pi$ projects the noise back into the $\ell_{\infty}$ ball with radius $\epsilon$, we initialize $\delta$ with Gaussian noise $\mathcal{N}(0,\epsilon^{2})$. After obtaining $\tilde{\mathbf{E}}$, we compute the adversarial loss as $)\mathcal{L}_{\text{adv}}=\mathcal{L_{\text{cls}}}(f(\tilde{\mathbf{E}}$)).
@@ 4.2 Consistency Regularization under Perturbations
While perturbation-based data augmentation exposes the model to diverse interventions, it does not explicitly constrain how the model’s intermediate representations ($Z_{I}$, $Z_{T}$) should respond to such perturbations. In multimodal settings, even small interventions on the task-irrelevant modality can undesirably shift the fused representations and lead to inconsistent predictions. To address this, we introduce a consistency regularization strategy that enforces output stability between original and perturbed inputs, serving as an indirect but effective constraint on the representations $Z_{I}$ and $Z_{T}$ to mitigate modality interference. By minimizing the divergence between the prediction distributions of original and perturbed inputs, the model is encouraged to maintain invariant behavior along the task-relevant causal paths. Specifically, to illustrate the role of consistency regularization, consider an image-heavy sample $(x_{I},x_{T},a)\in\mathcal{B}^{\text{img}}_{\text{orig}}$ and its perturbed version $(x_{I},\tilde{x}_{T},a)\in\mathcal{B}^{\text{img}}_{\text{pert}}$. The model’s predictive distribution over the answer $a$ is denoted as $p_{\theta}(a|x_{I},x_{T})$, parameterized by $\theta$. We define the consistency loss between the original and perturbed inputs as:
$\mathcal{L}_{\text{consistency}}=\mathrm{Consistency}\left(p_{\theta}(A|x_{I},x_{T})\,\|\,p_{\theta}(A|x_{I},\tilde{x}_{T})\right)$ (8)
This objective encourages the model to rely primarily on the visual representation $Z_{I}$ and to be invariant to perturbations in $X_{T}$, thus regularizing the causal path $Z_{I}\rightarrow A$ against interference. In practice, we calculate adopt multiple methods (e.g. KL, JS) based on the distribution of the output logit of the model produced by its language model head. Let $l^{\text{orig}}\in\mathbb{R}^{L_{A}\times V}$ and $l^{\text{pert}}\in\mathbb{R}^{L_{A}\times V}$ denote the pre-softmax logits produced by the language model for the original and perturbed samples respectively, where $L_{A}$ is the number of answer tokens and $V$ is the vocabulary size. We define the token-level consistency loss with temperature $\tau$ as (using KL divergence for illustration):
$\mathcal{L}_{\text{consistency}}=\frac{1}{L_{A}}\sum_{i=1}^{L_{A}}\sum_{v=1}^{V}\text{softmax}\left(\frac{l^{\text{orig}}_{i}}{\tau}\right)_{v}\cdot\log\left(\frac{\text{softmax}\left(\frac{l^{\text{orig}}_{i}}{\tau}\right)_{v}}{\text{softmax}\left(\frac{l^{\text{pert}}_{i}}{\tau}\right)_{v}}\right).$ (9)
The same principle applies in text-heavy tasks, where perturbations are instead applied to the image input $X_{I}$, and consistency is enforced between $p_{\theta}(A\mid x_{I},x_{T})$ and $p_{\theta}(A\mid\tilde{x}_{I},x_{T})$.
@@ 4.3 Final Training Objective
Our final training objective integrates both perturbation-based data augmentation and consistency regularization into a unified framework. For each batch, we begin with a set of original samples $\mathcal{B}_{\text{orig}}$ and dynamically construct their heuristic perturbed counterparts $\mathcal{B}_{\text{pert}}$ via input-level augmentations. We then apply PGD-based adversarial perturbations on both $\mathcal{B}_{\text{orig}}$ and $\mathcal{B}_{\text{pert}}$, and enforce consistency between the predictions of original and all perturbed samples. Specifically, we define the overall loss:
@@
5 Experiments
Models. We conduct experiments on three MLLM familiies with different parameter size: Qwen2.5-vl-3b qwen25vl , LLaVA-1.5-7B & LLaVA-1.5-13B llava15 and InstructBLIP-Vicuna-7B instructblip . These models typically integrate a pre-trained vision encoder (e.g., CLIP ViT-L/14 clip ) with a language model backbone (e.g., Vicuna vicuna ), connected via a modality projection module. Following the original LLaVA and Qwen-VL implementation, we freeze the vision encoder and train the multimodal projector and the language model. In our adaptation of InstructBLIP, we adopt an alternative tuning strategy: we freeze both the vision encoder and the Q-Former, and fine-tune the language model and the linear projection layer. We provide further discussion in  Appendix˜B.
Baselines. We include following baselines for comparison: LLaVA-1.5-13B + I-MoF mof : By applying the designed Interleaved Mixture-of-Features (I-MoF) module on LLaVA-1.5-13B llava15 to spatially combine CLIP clip and DINOv2 dinov2 visual tokens, it enhances visual grounding by integrating complementary features from contrastive and self-supervised vision encoders. VLMClassifier imagewikiqa : it enhances visually-grounded language models for image classification by fine-tuning them on ImageNet imagenet (VLMClassifier-1) or ImageNet combining LLaVA-Instruct llava (VLMClassifier-2).
Datasets. We evaluate models on benchmarks covering three task types: (i) Image-heavy tasks: Mini-ImageNet miniimagenet and Caltech-101 caltech101 , used in both training and evaluation, originally designed for image classification; (ii) Text-heavy tasks: OpenBookQA openbookqa and MMLU mmlu , consisting purely of textual question answering data; (iii) VQA tasks: For training, we use LLaVA-Instruct-dataset llava as the instruction-tuning dataset for related models. For InstructBLIP, we additionally include TextCaps textcaps , since these are the only publicly accessible VQA datasets known to be used in instruction-tuning for related models. For Qwen2.5-VL, whose instruction-tuning data is in-house and not publicly available, we use LLaVA-Instruct-dataset as a standardized alternative. For evaluation, we adopt three multiple-choice VQA benchmarks: ScienceQA-IMG scienceqa , MM-Bench-EN mmbench , and Seed-Bench-IMG seedbench . For ScienceQA and Seed-Bench, we only include examples with image context. For MM-Bench, we use the English version. All datasets are converted into a unified multiple-choice VQA format, enabling consistent modeling and evaluation across tasks and models. We report the accuracy of all multiple choice tasks and quantify the causal effect with the prediction change rate $\delta_{\text{cp}}$. All models are fine-tuned for 1 epoch with a fixed batch size $N_{\text{batch}}$. All the hyperparameters are listed in the  Appendix˜D.
Table 1: Evaluation of unimodal and multimodal tasks across multiple methods. For unimodal datasets, we report accuracy on origin setting (Orig) and the worst-performing perturbation (Perturbed). For VQA datasets, we report accuracy on the original setting. The best accuracy is marked in bold.
@@ 5 Experiments
   + VLMClassifier-1 imagewikiqa 15.1 0.0 15.6 0.0 61.5 61.5 47.8 47.5 61.1 36.2 35.9
   + VLMClassifier-2 imagewikiqa 15.6 0.0 15.1 0.0 61.8 61.2 47.4 47.8 61.8 35.8 36.0
@@
As shown in Table˜1, our method outperforms both VLMClassifier and I-MoF across different base MLLMs, demonstrating stronger robustness to modality interference and improved cross-modality competency. While I-MoF enhances visual grounding by integrating multiple visual features, it still suffers from modality interference: e.g. LLaVA-1.5-13B + I-MoF achieves 93.9% on original Mini-ImageNet but drops to 70.1% under perturbation (↓23.8%), indicating reliance on spurious textual cues. In contrast, our method maintains perturbed performance at 98.4% (↓0.1%). On the other hand, VLMClassifier, adopts vision-only fine-tuning, which leads to two critical limitations: vulnerability to cross-modal interference and degradation on VQA tasks, as LLaVA-1.5-7b + VLMClassifier-1 only reaches 35.8%/36.2% on MM-Bench/SeedBench, notably lower than both base LLaVA (72.1%/64.5%) and our method (73.7%/68.4%). These results highlight that vision-centric strategies, without addressing modality alignment, are insufficient for robust multimodal understanding. In text-heavy tasks such as OpenBookQA and MMLU, our method also achieves superior perturbed performance—82.1% vs. 64.5% on OpenBookQA, and 55.8% vs. 39.1% on MMLU (LLaVA-1.5-13B)—highlighting that addressing modality interference directly, rather than merely improving representations, is key to robust multimodal reasoning. Overall, unlike prior methods that often trade off between unimodal and multimodal performance, our method consistently improves both, achieving Pareto-optimality.
Table 2: Detailed multimodal reasoning accuracy (%) on multiple-choice VQA datasets across different ablation study settings. The best accuracy is marked in bold. Arrows (↑ / ↓) indicate relative changes compared to the Vanilla baseline of each model. Overall performance is computed as a weighted average across datasets, with weights proportional to each dataset’s test size.
@@
Table 3: Evaluation of Caltech-101 (Image-heavy) and MMLU (Text-heavy) across different ablation study settings. we report accuracy on the original setting (Orig) and the worst-performing perturbation (Perturbed).
Model Method Caltech-101 MMLU
Orig Perturbed Orig Perturbed
LLaVA-1.5-7B Vanilla 97.0 57.4 46.3 45.2
$D^{\text{VQA}}$ 96.2 46.3 46.8 45.6
$D^{\text{AUG}}$ 98.5 98.6 51.1 50.7
   +KL 98.6 98.7 51.1 50.7
   +PGD 98.7 98.5 50.6 50.3
Ours 99.3 99.0 51.5 51.0
InstructBLIP-7B Vanilla 90.3 17.5 35.3 35.2
$D^{\text{VQA}}$ 92.1 23.1 40.9 40.2
$D^{\text{AUG}}$ 99.0 56.1 50.0 49.7
   +JS 98.9 98.4 50.7 49.3
   +PGD 99.1 85.2 49.3 48.4
Ours 99.2 98.3 50.2 49.7
@@
To evaluate the effectiveness of each component in our framework, we conduct a comprehensive ablation study across multiple models and scales. We compare the pretrained models with the following strategies: FFT with $D^{\text{VQA}}$ (standard finetuning on VQA data), FFT with $D^{\text{AUG}}$ (supervised finetuning on mixed multi-task datasets with heuristic perturbations), FFT+KL/JS (adding consistency regularization on KL or JS divergence), FFT+RG (injecting random Gaussian noise instead of adversarial perturbation), FFT+PGD (FFT with heuristic perturbations and adversarial perturbations), and Ours (combining both perturbation-based data augmentation and consistency regularization).  Table˜2 presents overall VQA performance across different models and training strategies, while, Table˜3 evaluates model robustness under unimodal settings by comparing accuracy on the original inputs versus the worst-performing perturbation cases. Together, these results show the effectiveness of our method in improving both general VQA accuracy and robustness under modality interference. Table˜4 reports results on unimodal tasks with all the perturbations.  Table˜5 is full version of Table˜3. Across all model families (Qwen2.5-VL, InstructBLIP, and LLaVA-1.5) and model sizes (3B/7B/13B), our method consistently achieves best overall performance, improving accuracy on both unimodal and multimodal benchmarks. For instance, it boosts overall VQA accuracy (e.g., +4.7% on LLaVA-1.5-7B, +14.9% on InstructBLIP-7B), but also enhances robustness to modality interference–improve the performance under perturbations by over 50% on image-heavy tasks(e.g. 17.5% $\rightarrow$ 98.3% on InstructBLIP-7B on Caltech101).  Figure˜4 are radar chart visualizations. ( Table˜4,  Table˜5,  Figure˜4 in  Appendix˜C)
@@
We observe consistent improvements across both unimodal and multimodal tasks when moving from FFT w/ $D^{\text{VQA}}$ to FFT w/ $D^{\text{AUG}}$, highlighting the importance of incorporating modality-specific supervision and heuristic perturbations. For instance, on LLaVA-1.5-7B, both VQA accuracy (64.0% $\rightarrow$ 65.6%) and unimodal robustness (46.3% $\rightarrow$ 98.6% on Caltech101, 45.6% $\rightarrow$ 50.7% on MMLU) improve, indicating enhanced unimodal alignment. Building upon this, adding consistency regularization yields further gains by stabilizing model predictions under controlled perturbations on $X_{I}$ or $X_{T}$. Both KL and JS objectives lead to similar improvements, suggesting that the model equally benefits from all heuristic perturbations regardless of anchor choice. Finally, PGD-based adversarial perturbations outperform random Gaussian noise across almost all models, confirming that structured, learnable adversarial perturbation is more effective at weakening spurious modality shortcuts and encourages learning of more robust, task-relevant representations.

### v3

@@ 4 Methods
To mitigate modality interference and enhance cross-modality competency, we propose a unified perturbation-aware training framework that introduces interventions at both the input level (on $X_{I}$ and $X_{T}$) and the representation level (on $Z_{I}$ and $Z_{T}$) with consistency regularization. Overall pipeline is displayed in  Figure 3.
@@ 4.1 Perturbation-based Data Augmentation
To increase causal robustness along the desired paths $X_{I}\rightarrow Z_{I}\rightarrow A$ (for image-heavy tasks) or $X_{T}\rightarrow Z_{T}\rightarrow A$ (for text-heavy tasks), we first adopt a causally grounded data augmentation method by augmenting each sample with both heuristic perturbations and training time adversarial perturbations.
@@
We sample $N_{\text{img}}$ and $N_{\text{text}}$ examples from $\mathcal{D}^{\text{img}}$ and $\mathcal{D}^{\text{text}}$ to construct $\mathcal{B}^{\text{img}}_{\text{orig}}$ and $\mathcal{B}^{\text{text}}_{\text{orig}}$ respectively, and the remaining $N_{\text{vqa}}$ examples are VQA samples from $\mathcal{D}^{\text{vqa}}$ to construct $\mathcal{B}^{\text{vqa}}$ . With dynamically generated perturbed variants for each sample, the final training batch is:
@@
While heuristic perturbations simulate realistic but limited modality noise at the input level, they may not fully capture the worst-case failure modes of MLLMs, especially under complex spurious alignments in the representation space. To overcome this limitation, we introduce a stronger and more generalizable intervention through adversarial training. These perturbations simulate worst-case alignment disruptions during training, serving as targeted interventions on latent nodes $(Z_{I},Z_{T})$ to reveal the Direct Causal Effect (DCE) of irrelevant modalities on $A$. By optimizing the model under such adversarial conditions, we reduce the model’s reliance on spurious cross-modal signals and reinforce task-relevant causal pathways. Inspired by PGD (Madry et al., 2018; Chen et al., 2023b), we design a tailored perturbation strategy for multimodal token embeddings $(Z_{I},Z_{T})$. Unlike standard PGD that applies coarse sign-based updates, our method introduces two critical modifications: (1) Modality-specific perturbation masking, which restricts perturbations to task-irrelevant modalities via a binary mask, thereby transforming noise into targeted causal probes rather than indiscriminate corruption. (2) Raw-gradient updates, where we remove the sign operator and apply the raw gradient directly, yielding smoother, more diverse, and more realistic perturbations that better simulate modality interference. Formally, we construct perturbations $\delta=(\delta_{I},\delta_{T})$ in the latent space that maximize the model’s predictive loss:
@@
where $\epsilon$ bounds the perturbation strength and $f$ is the prediction function. We optimize $\delta$ through $n$ raw-gradient steps, updating at each step $t$ as:
$\delta^{(t+1)}=\Pi_{\|\delta\|_{\infty}\leq\epsilon}\Big(\delta^{(t)}+\alpha\cdot\nabla_{\delta}\mathcal{L_{\text{cls}}}(f(Z+\delta^{(t)}))\Big),$ (5)
where $\alpha$ is the step size, $\Pi$ projects the noise into the $\ell_{\infty}$ ball, and $Z=[Z_{I};Z_{T}]$ is the concatenated embedding. In practice, we integrate a modality-specific binary mask $M\in\{0,1\}^{L\times d}$, where $L$ is the sequence length and $d$ the hidden dimension, ensuring that perturbations only affect task-irrelevant tokens. Given multimodal embeddings $\mathbf{E}\in\mathbb{R}^{L\times d}$, the perturbed embeddings are:
@@
with $\odot$ denoting element-wise masking. We initialize $\delta$ with Gaussian noise $\mathcal{N}(0,\epsilon^{2})$ and update it for $T$ steps. The final adversarial objective is $)\mathcal{L}_{\text{adv}}=\mathcal{L}_{\text{cls}}(f(\tilde{\mathbf{E}}$)).
@@ 4.2 Consistency Regularization under Perturbations
[…]
$).\mathcal{L}_{\text{consistency}}=\mathrm{Consistency}\big(p_{\theta}(A|x)\,\|\,p_{\theta}(A|\tilde{x}$$\big)$. (7)
In practice, we instantiate this by applying distributional divergence (e.g., KL or JS) at the token level. Let $l^{\text{orig}},l^{\text{pert}}\in\mathbb{R}^{L_{A}\times V}$ denote the pre-softmax logits for the original and perturbed samples, where $L_{A}$ is the number of answer tokens and $V$ the vocabulary size. Using KL divergence with temperature $\tau$ as an example, the loss becomes (equally apply to image-heavy and text-heavy tasks):
$\mathcal{L}_{\text{consistency}}=\frac{1}{L_{A}}\sum_{i=1}^{L_{A}}\sum_{v=1}^{V}\text{softmax}\!\left(\tfrac{l^{\text{orig}}_{i}}{\tau}\right)_{v}\cdot\log\frac{\text{softmax}\!\left(\tfrac{l^{\text{orig}}_{i}}{\tau}\right)_{v}}{\text{softmax}\!\left(\tfrac{l^{\text{pert}}_{i}}{\tau}\right)_{v}}.$ (8)
@@ 4.3 Final Training Objective
Our final training objective integrates both perturbation-based data augmentation and consistency regularization into a unified framework. For each batch, we begin with a set of original samples $\mathcal{B}_{\text{orig}}$ and dynamically construct their heuristic perturbed counterparts $\mathcal{B}_{\text{pert}}$ via input-level augmentations. We then apply adversarial perturbations on both $\mathcal{B}_{\text{orig}}$ and $\mathcal{B}_{\text{pert}}$, and enforce consistency between the predictions of original and all perturbed samples. The overall loss is:
@@
Table 1: Evaluation on unimodal and VQA datasets. For unimodal datasets, we report accuracy under the original input (Orig) and the worst-performing perturbation (Perturbed). For VQA datasets, we report accuracy on the original setting. Best results are highlighted in bold.
@@
   + CoT (Wei et al., 2022) 81.7 28.9 80.9 36.0 38.8 38.9 39.6 38.7 64.7 65.2 64.1
   + VLMClassifier-1 (Zhang et al., 2024) 15.1 0.0 15.6 0.0 61.5 61.5 47.8 47.5 61.1 36.2 35.9
   + VLMClassifier-2 (Zhang et al., 2024) 15.6 0.0 15.1 0.0 61.8 61.2 47.4 47.8 61.8 35.8 36.0
@@
5 Experiments
Models. We conduct experiments on three MLLM families with different parameter size: Qwen2.5-vl-3b (Bai et al., 2025), LLaVA-1.5-7B & LLaVA-1.5-13B (Liu et al., 2023a) and InstructBLIP-Vicuna-7B (Luo et al., 2023). Following LLaVA and Qwen-VL, we freeze the vision encoder and train the multimodal projector and language model; for InstructBLIP, we instead freeze both the vision encoder and Q-Former, fine-tuning only the projection layer and language model (see Appendix B).
Baselines. We include following baselines for comparison: LLaVA-1.5-13B + I-MoF (Tong et al., 2024): By applying the designed Interleaved Mixture-of-Features (I-MoF) module on LLaVA-1.5-13B to spatially combine CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2023) visual tokens, it enhances visual grounding by integrating complementary features from contrastive and self-supervised vision encoders. VLMClassifier (Zhang et al., 2024): it enhances visually-grounded language models for image classification by fine-tuning them on ImageNet (Deng et al., 2009) (VLMClassifier-1) or ImageNet combining LLaVA-Instruct (Liu et al., 2023b) (VLMClassifier-2). Chain-of-Thought (CoT) Prompting (Wei et al., 2022): we further evaluate prompt-based mitigation by encouraging structured reasoning through CoT-style prompting. The specific prompt design and results are reported in Section C.4.
Datasets. We evaluate models on benchmarks covering three task types: (i) Image-heavy tasks: Mini-ImageNet (Russakovsky et al., 2015) and Caltech-101 (Fei-Fei et al., 2004), used in both training and evaluation, originally designed for image classification; (ii) Text-heavy tasks: OpenBookQA (Mihaylov et al., 2018) and MMLU (Hendrycks et al., 2020), consisting purely of textual question answering data; (iii) VQA tasks: For training, we use LLaVA-Instruct-dataset (Liu et al., 2023b) as the instruction-tuning dataset for related models. For InstructBLIP, we additionally use TextCaps (Sidorov et al., 2020) as another publicly available VQA dataset used in instruction tuning. For Qwen2.5-VL, whose instruction-tuning data is proprietary, we adopt LLaVA-Instruct as a standardized alternative. For evaluation, we adopt three multiple-choice VQA benchmarks: ScienceQA-IMG (Lu et al., 2022), MM-Bench-EN (Liu et al., 2023c), and Seed-Bench-IMG (Li et al., 2023). For ScienceQA and Seed-Bench, we only include examples with image context. For MM-Bench, we use the English version. All datasets are converted into a unified multiple-choice VQA format, enabling consistent modeling and evaluation across tasks and models. We report the accuracy of all multiple choice tasks and quantify the causal effect with the prediction change rate $\delta_{\text{cp}}$. All models are fine-tuned for 1 epoch with a fixed batch size $N_{\text{batch}}$. All the hyperparameters are listed in Appendix D. For each dataset, results are averaged over multiple independent runs. Following standard practice, inference is performed with deterministic decoding (temperature fixed at 0).
@@ 5 Experiments
As shown in Table 1, our method outperforms all baselines across different base MLLMs, demonstrating stronger robustness to modality interference and improved cross-modality competency. While CoT slightly improves performance in certain VQA settings, its overall gains are minimal and inconsistent, and it fails to mitigate modality interference under perturbed conditions (e.g., 85.9% vs. 97.9% on Mini-ImageNet). While I-MoF enhances visual grounding by integrating multiple visual features, it still suffers from modality interference: e.g. LLaVA-1.5-13B + I-MoF achieves 93.9% on original Mini-ImageNet but drops to 70.1% under perturbation (↓23.8%), indicating reliance on spurious textual cues. In contrast, our method maintains perturbed performance at 98.4% (↓0.1%). On the other hand, VLMClassifier, adopts vision-only fine-tuning, which leads to two critical limitations: vulnerability to cross-modal interference and degradation on VQA tasks, as LLaVA-1.5-7b + VLMClassifier-1 only reaches 35.8%/36.2% on MM-Bench/SeedBench, notably lower than both base LLaVA and our method (73.7%/68.4%). These results highlight that vision-centric strategies, without addressing modality alignment, are insufficient for robust multimodal understanding. In text-heavy tasks such as OpenBookQA and MMLU, our method also achieves superior perturbed performance(e.g., 55.8% vs. 39.1% on MMLU on LLaVA-1.5-13B)—highlighting that addressing modality interference directly, rather than merely improving representations, is key to robust multimodal reasoning. Overall, unlike prior methods that often trade off between unimodal and multimodal performance, our method consistently improves both, achieving Pareto-optimality.
Table 2: Multimodal reasoning accuracy (%) on VQA benchmarks under various ablation study configurations. The best accuracy is marked in bold. Arrows (↑ / ↓) indicate relative changes compared to the Vanilla baseline of each model. Overall performance is computed as a weighted average across datasets, with weights proportional to each dataset’s test size Table 13.(full results in Table 7)
@@
Table 3: Evaluation of Caltech-101 (image-heavy) and MMLU (text-heavy) across different ablation study settings (full results in Table 6). Model Method Caltech-101 MMLU Orig Perturbed Orig Perturbed LLaVA-1.5-7B Vanilla 97.0 57.4 46.3 45.2 $D^{\text{VQA}}$ 96.2 46.3 46.8 45.6 $D^{\text{AUG}}$ 98.5 98.6 51.1 50.7    +KL 98.6 98.7 51.1 50.7    +ADV 98.7 98.5 50.6 50.3 Ours 99.3 99.0 51.5 51.0 InstructBLIP-7B Vanilla 90.3 17.5 35.3 35.2 $D^{\text{VQA}}$ 92.1 23.1 40.9 40.2 $D^{\text{AUG}}$ 99.0 56.1 50.0 49.7    +JS 98.9 98.4 50.7 49.3    +ADV 99.1 85.2 49.3 48.4 Ours 99.2 98.3 50.2 49.7 Table 4: OOD robustness evaluation on Caltech-101 and MMLU, with OCR noise and Screenshot distractors (full results in Table 11). Model Method Caltech-101 MMLU Orig OCR Orig Screenshot LLaVA-1.5-7B Vanilla 97.0 92.8 46.3 44.8 $D^{\text{AUG}}$ 98.5 98.0 51.1 50.6    +ADV 98.7 98.4 50.6 51.3 Ours 99.3 99.0 51.5 51.3 InstructBLIP-7B Vanilla 90.3 83.6 35.3 34.9 $D^{\text{AUG}}$ 99.0 98.4 50.0 49.2    +ADV 99.1 98.8 49.3 49.3 Ours 99.2 99.0 50.2 49.2 Qwen2.5-VL-7B Vanilla 99.1 99.2 69.3 57.0 $D^{\text{AUG}}$ 99.7 99.3 70.4 63.7    +ADV 99.6 99.6 70.4 69.7 Ours 99.6 99.5 69.8 69.7
@@
To evaluate the effectiveness of each component in our framework, we conduct a comprehensive ablation study across multiple models and scales. We compare the pretrained models with the following strategies: FFT with $D^{\text{VQA}}$ (standard finetuning on VQA data), FFT with $D^{\text{AUG}}$ (supervised finetuning on mixed multi-task datasets with heuristic perturbations), FFT+KL/JS (adding consistency regularization on KL or JS divergence), FFT+RG (injecting random Gaussian noise into token embeddings), FFT+ADV (FFT with heuristic & adversarial perturbations), and Ours (combining both perturbation-based data augmentation and consistency regularization).  Table 2 presents overall VQA performance, and  Table 3 evaluates model robustness under unimodal settings. ( Table 5 reports results with all perturbations.) Together, these results show the effectiveness of our method in improving both general VQA accuracy and robustness under modality interference. Across all model families (Qwen2.5-VL, InstructBLIP, LLaVA-1.5) and model sizes (3B/7B/13B), our method consistently achieves best overall performance, improving accuracy on both unimodal and multimodal benchmarks. For instance, it boosts overall VQA accuracy (e.g., +14.9% on InstructBLIP-7B), but also enhances robustness to modality interference–improve the performance under perturbations by over 50% on image-heavy tasks(e.g. 17.5% $\rightarrow$ 98.3% with InstructBLIP-7B on Caltech101). We also extend evaluation from MCQA to free-form QA (Section C.5).
@@
We observe consistent improvements across both unimodal and multimodal tasks when moving from FFT w/ $D^{\text{VQA}}$ to FFT w/ $D^{\text{AUG}}$, highlighting the importance of incorporating modality-specific supervision and heuristic perturbations. Building upon this, adding consistency regularization yields further gains by stabilizing model predictions under controlled perturbations on $X_{I}$ or $X_{T}$. Both KL and JS objectives lead to similar improvements, suggesting that the model equally benefits from all heuristic perturbations regardless of anchor choice.Finally, we compare adversarial perturbations with random Gaussian noise, and find that FFT+ADV consistently outperforms FFT+RG across most backbones, indicating that structured perturbations more effectively suppress spurious shortcuts and promote robust, task-relevant representations. To further validate generalization, we introduce two real-world out-of-distribution perturbations at test time: (i) noisy OCR snippets sampled from FUNSD (Jaume et al., 2019) as irrelevant text into image-heavy tasks; and (ii) unrelated UI screenshots from RICO (Deka et al., 2017) as distractor images in text-heavy tasks. As shown in Table 4, adversarial training significantly improves robustness under these unseen perturbations with consistent gains (e.g., on InstructBLIP-7B, 83.6% $\rightarrow$ 99.0% under OCR noise). These results demonstrate that the modest overhead of adversarial training (see Appendix E) yields substantial gains in out-of-domain generalization, a crucial property for reliable deployment.